# On estimation and order selection for multivariate extremes via clustering

### v2

@@ 2.3 Spherical clustering for multivariate extremes
Now we consider applying the $k$-clustering in Definition 2.3 to the random subsample ${W}_{n}$. In view of Proposition 2.5, the distribution of $W_{n}$ on $\mathbb{S}_{+}^{d-1}$ serves as a good approximation to $H$. When $H$ is a discrete measure with finitely many atoms, clustering can be applied to $W_{n}$ in such a way that the cluster centers closely approximate the locations of the atoms, and the cluster proportions provide accurate estimates of the associated probabilities. Furthermore, these estimators are shown to be consistent, as demonstrated in the following Corollary.
Corollary 2.6.
Suppose $\mathbf{X}$ is as in Proposition 2.5, and furthermore, has a spectral measure of the following form:
@@
where $\mathbf{a}_{i}$’s are distinct points on $\mathbb{S}_{+}^{d-1}$, and $p_{i}>0$, $p_{1}+\ldots+p_{k}=1$. Let $W_{n}$ denote the extremal subsample as in (14), and $\left({A}_{k,n}=\left({\mathbf{a}}_{1,n}^{k},\ldots,{\mathbf{a}}_{k,n}^{k}\right),\mathfrak{C}_{k,n}=\left\{{C}_{1,n}^{k},\ldots,{C}_{k,n}^{k}\right\}\right)$ form a $k$-clustering of $W_{n}$ as defined in Definition 2.3 with respect to a dissimilarity measure $D$ defined in Definition 2.1. Let
@@
The convergence of $\mathbf{a}_{\pi_{n}(i),n}^{k}$ follows from [23, Theorem 3.1] (stated as convergence in Hausdorff distance between ${A}_{k,n}$ and $\{\mathbf{a}_{1},\ldots,\mathbf{a}_{k}\}$) and Proposition 2.5 above; see also the discussion in [23, Section 4]). It remains to show the convergence of $p_{\pi_{n}(i),n}^{k}$, $i\in\{1,\ldots,k\}$. Set
@@
Fix $\epsilon\in(0,r_{A}/3)$. By what has been proved and the continuity of $D^{\dagger}$, at almost every outcome $\omega$ of the sample space $\Omega$, for $n$ sufficiently large, we have the dual dissimilarity $D^{\dagger}(\mathbf{a}_{\pi_{n}(i),n}^{k},\mathbf{a}_{i})<\epsilon$ , $i\in\{1,\ldots,k\}$. Fix for now such an $\omega$ and let $n$ be sufficiently large (possibly depending on $\omega$). Then by the triangular inequality (10), we have $B_{D}(\mathbf{a}_{i},\epsilon)\subset B_{D}\left(\mathbf{a}_{\pi_{n}(i),n}^{k},2\epsilon\right)\subset B_{D}(\mathbf{a}_{i},3\epsilon)$, $i\in\{1,\ldots,k\}$. Note that $B_{D}\left(\mathbf{a}_{\pi_{n}(i),n}^{k},2\epsilon\right)\cap W_{n}$ are disjoint for $i\in\{1,\ldots,k\}$ due to the choice of $\epsilon$. So in view of Definition 2.3, we have $B_{D}(\mathbf{a}_{i},\epsilon)\cap W_{n}\subset B_{D}\left(\mathbf{a}_{\pi_{n}(i),n}^{k},2\epsilon\right)\cap W_{n}\subset C_{\pi_{n}(i),n}^{k}\subset\left(B_{D}(\mathbf{a}_{i},\epsilon)\cup\cap_{j\neq i}B_{D}(\mathbf{a}_{j},\epsilon)^{c}\right)\cap W_{n}$, and hence
@@
Recall the inequality above is derived for almost every $\omega$ with $n$ sufficiently large. The conclusion then follows from Proposition 2.5 since both sides above converges almost surely to $p_{i}$ as $n\rightarrow\infty$. ∎
Remark 2.7.
Comparing [23, Proposition 3.3] with Proposition 2.5 and Corollary 2.6 here, we have chosen to work directly under the marginal standardization assumption in (1) and not to treat the empirical marginal transformations as in [23, Eq. (3.5)] for simplicity. Nevertheless, the consistency result of the order selection below (Theorem 3.1) can be extended to the setup of [23] based on the results there.
@@ 3.1 The method
Following the notation and setup in Section 2.2, suppose $W$ is a multiset on $\mathbb{S}_{+}^{d-1}$ and $1\leq k\leq|W|<\infty$. Let $\left(A_{k}^{*}=\left\{\mathbf{a}_{1}^{*},\ldots,\mathbf{a}_{k}^{*}\right\},\mathfrak{C}_{k}=\left\{C_{1},\ldots,C_{k}\right\}\right)$ be a $k$-clustering of $W$ with respect to a dissimilarity measure $D$ as in Definition 2.3. Define for $\mathbf{w}\in W$ that
@@
Motivated by these observations, we propose to introduce a penalty term that discourages small cluster size and small $D$ dissimilarity between centers. Recall that for a $k$-clustering $\left(A_{k}^{*}=\left\{\mathbf{a}_{1}^{*},\ldots,\mathbf{a}_{k}^{*}\right\},\mathfrak{C}_{k}=\left\{C_{1},\ldots,C_{k}\right\}\right)$ of a multiset $W$, the set $\left\{\mathbf{a}_{1}^{*},\ldots,\mathbf{a}_{k}^{*}\right\}$ represents the cluster centers and $\left\{C_{1},\ldots,C_{k}\right\}$ denotes the partition of $W$. Then $\min_{i=1,\ldots,k}\left(\frac{|C_{i}|}{|W|/k}\right)$ is the smallest cluster size relative to the average cluster size $|W|/k$, and $\min_{1\leq i<j\leq k}D\left(\mathbf{a}_{i}^{*},\mathbf{a}_{j}^{*}\right)$ is the smallest pairwise dissimilarity between centers. Note that both quantities are between 00 and $1$ (recall that $D$ maps onto $[0,1]$), and we want to introduce a criterion that penalizes either of them being small. There is arguably some arbitrariness in designing this penalty criterion. Through some mathematical heuristics and extensive experiments, we find that the following penalty works relatively well. Let $t\geq 0$ be a tuning parameter. Set
@@
Theorem 3.1.
Suppose $\mathbf{X}$ satisfying (1) and (2) has a discrete spectral measure of the form $H=\sum_{i=1}^{k}p_{i}\delta_{\mathbf{a}_{i}}$, $k\in\mathbb{Z}_{+}$ where $\mathbf{a}_{i}$’s are distinct points on $\mathbb{S}_{+}^{d-1}$, and $p_{i}>0$, $p_{1}+\ldots+p_{k}=1$. Let $W_{n}$ denote the extremal subsample as in (14) with $\ell_{n}\rightarrow\infty$, $\ell_{n}/n\rightarrow 0$, and $\left(A_{m,n},\mathfrak{C}_{m,n}\right)$, $m\in\mathbb{Z}_{+}$, form an $m$-clustering of $W_{n}$ as defined in Definition 2.3 with respect to a dissimilarity measure $D$ defined in Definition 2.1. Let $r_{A}$ be defined as in (20) and define
@@
The theorem implies that as long as the tuning parameter is in an appropriate range, with probability tending to $1$ as $n\rightarrow\infty$, the true order $m=k$ uniquely maximizes the penalized ASW. The proof of Theorem 3.1 can be found in Section 3.2. In Proposition 4.5 below, we will provide a rate of how fast the probability of false order selection decays to zero.
@@ 3.2 Consistency of order selection via penalized silhouette
In this section, we prove Theorem 3.1.
@@ 3.2.1 Some deterministic estimates
We first prepare some deterministic estimates regarding the $k$-clustering in Definition 2.3 and ASW in (21). Based on Proposition 2.5, if the true spectral measure $H$ consists of finitely many atoms, and we select the subset $S$ in the proposition as a union of neighborhoods around each atom, then, for sufficiently large $n$, almost all points in the extremal subsample $W_{n}$ in (14) will lie near one of these atoms with respect to the dissimilarity measure $D$. This motivates us to address a scenario where we aim to cluster a multiset where the majority of points are concentrated around a finite number of centers.
Fix throughout this section distinct points $\left\{\mathbf{a}_{1},\ldots,\mathbf{a}_{k}\right\}=:A$ on $\mathbb{S}_{+}^{d-1}$, $k\in\mathbb{Z}_{+}$, $p_{i}>0$ with $p_{1}+\ldots+p_{k}=1$, and a dissimilarity measure $D$ defined in Definition 2.1. We say a multiset $W$ on $\mathbb{S}_{+}^{d-1}$ with $|W|<\infty$ satisfies the concentration condition $\mathcal{A}(\epsilon,\delta)$ with $\epsilon,\delta>0$, if $k\leq|W|\wedge\delta^{-1}$ such that $\epsilon\in(0,r_{A})$ with $r_{A}$ in (20), $\delta\in\left(0,p_{\min}\right)$ with $p_{\min}$ in (24), and
@@
Remark 3.2.
Note that $r_{A}^{\dagger}(s)>0$ for any $s>0$, and $r_{A}^{\dagger}(s)\rightarrow 0$ as $s\rightarrow 0$; see Remark 2.2. For sufficiently small $\epsilon$, we always have $r_{A}^{\dagger}(\epsilon)<r_{A}$.
@@
Lemma 3.3.
Suppose a multiset $W$ satisfies the condition $\mathcal{A}(\epsilon,\delta)$ and $1\leq m<k$. Let $(A^{*}_{m},\mathfrak{C}_{m})$ be an $m$-clustering of $W$ as defined in Definition 2.3. Then the (unpenalized) ASW $\bar{S}$ satisfies
@@
Lemma 3.4.
Suppose a multiset $W$ satisfies the condition $\mathcal{A}(\epsilon,\delta)$. Let $\left(A^{*}_{m}=\{\mathbf{a}_{1}^{*},\ldots,\mathbf{a}_{m}^{*}\},\mathfrak{C}_{m}\right)$ be an $m$-clustering of $W$ as defined in Definition 2.3, $m\geq k$. Then for any $i\in\{1,\ldots,k\}$, there exists $j\in\{1,\ldots,m\}$, such that $D\left(\mathbf{a}_{j}^{*},\mathbf{a}_{i}\right)<\epsilon^{\prime}$, where
@@
Lemma 3.5.
Suppose a multiset $W$ satisfies the condition $\mathcal{A}(\epsilon,\delta)$. Assume additionally that $\epsilon^{\prime}$ in (28) satisfies $\epsilon^{\prime}<r_{A}$. Let $\left(A^{*}_{m}=\{\mathbf{a}_{1}^{*},\ldots,\mathbf{a}_{m}^{*}\},\mathfrak{C}_{m}=\{C_{1},\ldots,C_{m}\}\right)$ be an $m$-clustering of $W$ as defined in Definition 2.3, $m>k$. Then either of the following happens:
@@
Proof.
Since $B_{D}\left(\mathbf{a}_{i},\epsilon^{\prime}\right)$, $i\in\left\{1,\ldots,k\right\}$, are disjoint (because $\epsilon^{\prime}<r_{A}$), by Lemma 3.4, we can, without loss of generality, assume that $\mathbf{a}_{i}^{*}\in B_{D}\left(\mathbf{a}_{i},\epsilon^{\prime}\right)$, $i\in\left\{1,\ldots,k\right\}$ We now divide into two cases as follows.
@@
This in view of Definition 2.3 implies that $W\cap B_{D}\left(\mathbf{a}_{i},\epsilon\right)\subset W\cap C_{j}^{c}$ for all $i\in\left\{1,\ldots,k\right\}$. Therefore, we have by (26) that
@@
Lemma 3.6.
Suppose a multiset $W$ satisfies the condition $\mathcal{A}(\epsilon,\delta)$. Let $\left(A^{*}_{k}=\{\mathbf{a}_{1}^{*},\ldots,\mathbf{a}_{k}^{*}\},\mathfrak{C}_{k}=\{C_{1},\ldots,C_{k}\}\right)$ be a $k$-clustering of $W$ as defined in Definition 2.3. Suppose in addition
@@
In addition, with the same permutation $\pi:\{1,\ldots,k\}\mapsto\{1,\ldots,k\}$ found in Lemma 3.4, we have
@@
Proof.
Since $r_{A}>\epsilon^{\prime}$, by Lemma 3.4, there exists a permutation $\pi:\{1,\ldots,k\}\mapsto\{1,\ldots,k\}$, such that $D\left(\mathbf{a}_{i},\mathbf{a}_{\pi(i)}^{*}\right)<\epsilon^{\prime}$, $i\in\left\{1,\ldots,k\right\}$. Then for each $i$ and any $\mathbf{w}\in B_{D}(\mathbf{a}_{i},\epsilon)$, we have by the triangular inequality (10) that
@@
For the second claim, in view of Definition 2.3, (31), (32), (33), we have $W\cap B_{D}\left(\mathbf{a}_{i},\epsilon\right)\subset C_{\pi(i)}$, $i\in\left\{1,\ldots,k\right\}$. Hence by (26),

### v3

@@ 2.3 Spherical clustering for multivariate extremes
Now we consider applying the $k$-clustering in Definition 2 to the random subsample ${W}_{n}$. In view of Proposition 1, the distribution of $W_{n}$ on $\mathbb{S}_{+}^{d-1}$ serves as a good approximation to $H$. When $H$ is a discrete measure with finitely many atoms, clustering can be applied to $W_{n}$ in such a way that the cluster centers closely approximate the locations of the atoms, and the cluster proportions provide accurate estimates of the associated probabilities. Furthermore, these estimators are shown to be consistent, as demonstrated in the following Corollary.
Corollary 1.
Suppose $\mathbf{X}$ is as in Proposition 1, and furthermore, has a spectral measure of the following form:
@@
where $\mathbf{a}_{i}$’s are distinct points on $\mathbb{S}_{+}^{d-1}$, and $p_{i}>0$, $p_{1}+\ldots+p_{k}=1$. Let $W_{n}$ denote the extremal subsample as in (14), and $\left({A}_{k,n}=\left({\mathbf{a}}_{1,n}^{k},\ldots,{\mathbf{a}}_{k,n}^{k}\right),\mathfrak{C}_{k,n}=\left\{{C}_{1,n}^{k},\ldots,{C}_{k,n}^{k}\right\}\right)$ form a $k$-clustering of $W_{n}$ as defined in Definition 2 with respect to a dissimilarity measure $D$ defined in Definition 1. Let
@@
The convergence of $\mathbf{a}_{\pi_{n}(i),n}^{k}$ follows from [23, Theorem 3.1] (stated as convergence in Hausdorff distance between ${A}_{k,n}$ and $\{\mathbf{a}_{1},\ldots,\mathbf{a}_{k}\}$) and Proposition 1 above; see also the discussion in [23, Section 4]). It remains to show the convergence of $p_{\pi_{n}(i),n}^{k}$, $i\in\{1,\ldots,k\}$. Set
@@
Fix $\epsilon\in(0,r_{A}/3)$. By what has been proved and the continuity of $D^{\dagger}$, at almost every outcome $\omega$ of the sample space $\Omega$, for $n$ sufficiently large, we have the dual dissimilarity $D^{\dagger}(\mathbf{a}_{\pi_{n}(i),n}^{k},\mathbf{a}_{i})<\epsilon$ , $i\in\{1,\ldots,k\}$. Fix for now such an $\omega$ and let $n$ be sufficiently large (possibly depending on $\omega$). Then by the triangular inequality (10), we have $B_{D}(\mathbf{a}_{i},\epsilon)\subset B_{D}\left(\mathbf{a}_{\pi_{n}(i),n}^{k},2\epsilon\right)\subset B_{D}(\mathbf{a}_{i},3\epsilon)$, $i\in\{1,\ldots,k\}$. Note that $B_{D}\left(\mathbf{a}_{\pi_{n}(i),n}^{k},2\epsilon\right)\cap W_{n}$ are disjoint for $i\in\{1,\ldots,k\}$ due to the choice of $\epsilon$. So in view of Definition 2, we have $B_{D}(\mathbf{a}_{i},\epsilon)\cap W_{n}\subset B_{D}\left(\mathbf{a}_{\pi_{n}(i),n}^{k},2\epsilon\right)\cap W_{n}\subset C_{\pi_{n}(i),n}^{k}\subset\left(B_{D}(\mathbf{a}_{i},\epsilon)\cup\cap_{j\neq i}B_{D}(\mathbf{a}_{j},\epsilon)^{c}\right)\cap W_{n}$, and hence
@@
Recall the inequality above is derived for almost every $\omega$ with $n$ sufficiently large. The conclusion then follows from Proposition 1 since both sides above converges almost surely to $p_{i}$ as $n\rightarrow\infty$. ∎
Remark 3.
Comparing [23, Proposition 3.3] with Proposition 1 and Corollary 1 here, we have chosen to work directly under the marginal standardization assumption in (1) and not to treat the empirical marginal transformations as in [23, Eq. (3.5)] for simplicity. Nevertheless, the consistency result of the order selection below (Theorem 1) can be extended to the setup of [23] based on the results there.
@@ 3.1 The method
Following the notation and setup in Section 2.2, suppose $W$ is a multiset on $\mathbb{S}_{+}^{d-1}$ and $1\leq k\leq|W|<\infty$. Let $\left(A_{k}^{*}=\left\{\mathbf{a}_{1}^{*},\ldots,\mathbf{a}_{k}^{*}\right\},\mathfrak{C}_{k}=\left\{C_{1},\ldots,C_{k}\right\}\right)$ be a $k$-clustering of $W$ with respect to a dissimilarity measure $D$ as in Definition 2. Define for $\mathbf{w}\in W$ that
@@
Motivated by these observations, we propose to introduce a penalty term that discourages small cluster size and small $D$ dissimilarity between centers. Recall that for a $k$-clustering $\left(A_{k}^{*}=\left\{\mathbf{a}_{1}^{*},\ldots,\mathbf{a}_{k}^{*}\right\},\mathfrak{C}_{k}=\left\{C_{1},\ldots,C_{k}\right\}\right)$ of a multiset $W$, the set $\left\{\mathbf{a}_{1}^{*},\ldots,\mathbf{a}_{k}^{*}\right\}$ represents the cluster centers and $\left\{C_{1},\ldots,C_{k}\right\}$ denotes the partition of $W$. Then $\min_{i=1,\ldots,k}\left(|C_{i}|/(|W|/k)\right)$ is the smallest cluster size relative to the average cluster size $|W|/k$, and $\min_{1\leq i<j\leq k}D\left(\mathbf{a}_{i}^{*},\mathbf{a}_{j}^{*}\right)$ is the smallest pairwise dissimilarity between centers. Note that both quantities are between 00 and $1$ (recall that $D$ maps onto $[0,1]$), and we want to introduce a criterion that penalizes either of them being small. There is arguably some arbitrariness in designing this penalty criterion. Through some mathematical heuristics and extensive experiments, we find that the following penalty works relatively well. Let $t\geq 0$ be a tuning parameter. Set
@@
Theorem 1.
Suppose $\mathbf{X}$ satisfying (1) and (2) has a discrete spectral measure of the form $H=\sum_{i=1}^{k}p_{i}\delta_{\mathbf{a}_{i}}$, $k\in\mathbb{Z}_{+}$ where $\mathbf{a}_{i}$’s are distinct points on $\mathbb{S}_{+}^{d-1}$, and $p_{i}>0$, $p_{1}+\ldots+p_{k}=1$. Let $W_{n}$ denote the extremal subsample as in (14) with $\ell_{n}\rightarrow\infty$, $\ell_{n}/n\rightarrow 0$, and $\left(A_{m,n},\mathfrak{C}_{m,n}\right)$, $m\in\mathbb{Z}_{+}$, form an $m$-clustering of $W_{n}$ as defined in Definition 2 with respect to a dissimilarity measure $D$ defined in Definition 1. Let $r_{A}$ be defined as in (20) and define
@@
The theorem implies that as long as the tuning parameter is in an appropriate range, with probability tending to $1$ as $n\rightarrow\infty$, the true order $m=k$ uniquely maximizes the penalized ASW. The proof of Theorem 1 can be found in Section 3.2. In Proposition 5 below, we will provide a rate of how fast the probability of false order selection decays to zero.
@@ 3.2 Consistency of order selection via penalized silhouette
In this section, we prove Theorem 1.
@@ 3.2.1 Some deterministic estimates
We first prepare some deterministic estimates regarding the $k$-clustering in Definition 2 and ASW in (21). Based on Proposition 1, if the true spectral measure $H$ consists of finitely many atoms, and we select the subset $S$ in the proposition as a union of neighborhoods around each atom, then, for sufficiently large $n$, almost all points in the extremal subsample $W_{n}$ in (14) will lie near one of these atoms with respect to the dissimilarity measure $D$. This motivates us to address a scenario where we aim to cluster a multiset where the majority of points are concentrated around a finite number of centers.
Fix throughout this section distinct points $\left\{\mathbf{a}_{1},\ldots,\mathbf{a}_{k}\right\}=:A$ on $\mathbb{S}_{+}^{d-1}$, $k\in\mathbb{Z}_{+}$, $p_{i}>0$ with $p_{1}+\ldots+p_{k}=1$, and a dissimilarity measure $D$ defined in Definition 1. We say a multiset $W$ on $\mathbb{S}_{+}^{d-1}$ with $|W|<\infty$ satisfies the concentration condition $\mathcal{A}(\epsilon,\delta)$ with $\epsilon,\delta>0$, if $k\leq|W|\wedge\delta^{-1}$ such that $\epsilon\in(0,r_{A})$ with $r_{A}$ in (20), $\delta\in\left(0,p_{\min}\right)$ with $p_{\min}$ in (24), and
@@
Remark 4.
Note that $r_{A}^{\dagger}(s)>0$ for any $s>0$, and $r_{A}^{\dagger}(s)\rightarrow 0$ as $s\rightarrow 0$; see Remark 1. For sufficiently small $\epsilon$, we always have $r_{A}^{\dagger}(\epsilon)<r_{A}$.
@@
Lemma 1.
Suppose a multiset $W$ satisfies the condition $\mathcal{A}(\epsilon,\delta)$ and $1\leq m<k$. Let $(A^{*}_{m},\mathfrak{C}_{m})$ be an $m$-clustering of $W$ as defined in Definition 2. Then the (unpenalized) ASW $\bar{S}$ satisfies
@@
Lemma 2.
Suppose a multiset $W$ satisfies the condition $\mathcal{A}(\epsilon,\delta)$. Let $\left(A^{*}_{m}=\{\mathbf{a}_{1}^{*},\ldots,\mathbf{a}_{m}^{*}\},\mathfrak{C}_{m}\right)$ be an $m$-clustering of $W$ as defined in Definition 2, $m\geq k$. Then for any $i\in\{1,\ldots,k\}$, there exists $j\in\{1,\ldots,m\}$, such that $D\left(\mathbf{a}_{j}^{*},\mathbf{a}_{i}\right)<\epsilon^{\prime}$, where
@@
Lemma 3.
Suppose a multiset $W$ satisfies the condition $\mathcal{A}(\epsilon,\delta)$. Assume additionally that $\epsilon^{\prime}$ in (28) satisfies $\epsilon^{\prime}<r_{A}$. Let $\left(A^{*}_{m}=\{\mathbf{a}_{1}^{*},\ldots,\mathbf{a}_{m}^{*}\},\mathfrak{C}_{m}=\{C_{1},\ldots,C_{m}\}\right)$ be an $m$-clustering of $W$ as defined in Definition 2, $m>k$. Then either of the following happens:
@@
Proof:.
Since $B_{D}\left(\mathbf{a}_{i},\epsilon^{\prime}\right)$, $i\in\left\{1,\ldots,k\right\}$, are disjoint (because $\epsilon^{\prime}<r_{A}$), by Lemma 2, we can, without loss of generality, assume that $\mathbf{a}_{i}^{*}\in B_{D}\left(\mathbf{a}_{i},\epsilon^{\prime}\right)$, $i\in\left\{1,\ldots,k\right\}$ We now divide into two cases as follows.
@@
This in view of Definition 2 implies that $W\cap B_{D}\left(\mathbf{a}_{i},\epsilon\right)\subset W\cap C_{j}^{c}$ for all $i\in\left\{1,\ldots,k\right\}$. Therefore, we have by (26) that
@@
Lemma 4.
Suppose a multiset $W$ satisfies the condition $\mathcal{A}(\epsilon,\delta)$. Let $\left(A^{*}_{k}=\{\mathbf{a}_{1}^{*},\ldots,\mathbf{a}_{k}^{*}\},\mathfrak{C}_{k}=\{C_{1},\ldots,C_{k}\}\right)$ be a $k$-clustering of $W$ as defined in Definition 2. Suppose in addition
@@
In addition, with the same permutation $\pi:\{1,\ldots,k\}\mapsto\{1,\ldots,k\}$ found in Lemma 2, we have
@@
Proof:.
Since $r_{A}>\epsilon^{\prime}$, by Lemma 2, there exists a permutation $\pi:\{1,\ldots,k\}\mapsto\{1,\ldots,k\}$, such that $D\left(\mathbf{a}_{i},\mathbf{a}_{\pi(i)}^{*}\right)<\epsilon^{\prime}$, $i\in\left\{1,\ldots,k\right\}$. Then for each $i$ and any $\mathbf{w}\in B_{D}(\mathbf{a}_{i},\epsilon)$, we have by the triangular inequality (10) that
@@
For the second claim, in view of Definition 2, (31), (32), (33), we have $W\cap B_{D}\left(\mathbf{a}_{i},\epsilon\right)\subset C_{\pi(i)}$, $i\in\left\{1,\ldots,k\right\}$. Hence by (26),